I need help with this question:
a plane flies 440 miles with the wind and 330 in the same length of time If the speed of the wind is 21. what is the speed of the plane in still​ air?

Answer by Boreal(15235)

I'm assuming the plane's speed x allows it to travel 440 miles with the wind and in the same time 330 miles against the wind.
(x+21)t=440, t=440/(x+21); rate x time =distance. time=distance/rate
(x-21)t=330, t=330/(x-21)
Set the two values equal to each other.
440/(x+21)=330/(x-21)
cross multiply
440x-9240=330x+6930
110x=16170
x=147 mph ANSWER
with the wind, 168 mph and 440/168=2.619 hours
against the wind, 126 mph and 330/126=2.619 hours, checks.